OpenStudy (anonymous):

integrate Cos5xCos3x

OpenStudy (anonymous):

ans = 0.5(sin8x/8+sin2x/2)

OpenStudy (anonymous):

+ c

OpenStudy (dumbcow):

using integration by parts i got something a little different \[\int\limits_{}^{} \cos 5x \cos 3x = -\frac{3}{16}\sin 3x \cos 5x + \frac{5}{16}\sin 5x \cos 3x +C\]